OpenStudy (anonymous):

The ordered pairs (1, 3), (2, 9), (3, 27), (4, 81), and (5, 243) represent a function. What is a rule that represents this function? y = x3 y = 3x y = 3x y = x + 2

OpenStudy (anonymous):

your 2nd and 3rd choices appear to be the same thing. Does one of them have an exponent?

OpenStudy (anonymous):

The 3rd one. the x is the exponent, sorry.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

y = 3^x \[y = 3^{x}\] But you should learn to check out the points in the rule to make sure the rule works.

OpenStudy (anonymous):

Better for learning than just having someone give you the answers...

OpenStudy (anonymous):

y = 3^x

OpenStudy (anonymous):

Notice that the first rule is obviously wrong... y = x^3 doesn't even give you the first point correctly... with x = 1, y = x^3 is still = 1, but so it doesn't work for the point (1, 3)

OpenStudy (anonymous):

y = 3x and y = x + 2 both work fine for the first point, but not for any of the rest of the points...

OpenStudy (anonymous):

Thank you for explaining it to me.

OpenStudy (anonymous):

when you notice the y-values getting really big quickly like this, that's often a hint that there is an exponent involved... having x in the exponent makes the y values get huge fast :)

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

Refer to the attachement.